OpenStudy (anonymous):

The length of a rectangle is twice the width. The area is 72 yd^2. Find the length and the width

OpenStudy (anonymous):

length = x width = 2x 2x(x)= 72 yd^2 2x^2= 72 x^2= 36 x=6

OpenStudy (anonymous):

so the length is going to be 36 and the width is going to be 6

OpenStudy (anonymous):

nope, width= 6 length= 12

OpenStudy (anonymous):

Find and label for Algebra the vertex and the line of symmetry for f(x)=2(x-3)^2 The vertex is (type an ordered pair for the vertex). The equation of the line of symmetry is x=